OpenStudy (anonymous):

just verify if I am right, please... INTEGRATION (haven't learned by parts method, don't think it is needed tho)

OpenStudy (anonymous):

\[\LARGE \int\limits_{ }^{ }\frac{\sin(2x)}{1+\cos^2(x)}~dx\]

OpenStudy (anonymous):

\[u=1+\cos^2(x)\]\[\frac{du}{dx}(1+\cos^2(x))=\frac{du}{dx}(\cos~x)^2=-2\cos(x)\sin(x)=-\sin(2x)\]\[-du=\sin(2x)~dx\]

OpenStudy (anonymous):

\[-\LARGE \int\limits_{ }^{ }\frac{1}{u}~dx\] \[answer,~~~\LARGE -\ln(1+\cos^2x)+C\]

OpenStudy (anonymous):

This is the first problem, and the second one is....

OpenStudy (anonymous):

\[\LARGE \int\limits_{ }^{ }\frac{\sin(x)}{1+\cos^2(x)}~dx\]This is where I am a little bit stuck

OpenStudy (anonymous):

Anyone that can help me? @mathstudent55 can you give me a suggestion?

OpenStudy (anonymous):

No, I got it.... \[u=\cos(x),~-du=\sin(x)\] \[-\LARGE \int\limits_{ }^{ }\frac{1}{1+u^2}~du\] -tan^(-1) (cos x) + C

OpenStudy (anonymous):

got it, no need helpin me,